# Unfolding ADMM for Enhanced Subspace Clustering of Hyperspectral Images thanks: This work was supported by the Research Foundation - Flanders, under the grant G094122N (project SPYDER), the Flanders AI Research Programme under Grant 174B09119, and by the China Scholarship Council under Grant 202106150007

### v2

@@ Abstract
Deep subspace clustering methods are now prominent in clustering, typically using fully connected networks and a self-representation loss function. However, these methods often struggle with overfitting and lack interpretability. In this paper, we explore an alternative clustering approach based on deep unfolding. By unfolding iterative optimization methods into neural networks, this approach offers enhanced interpretability and reliability compared to data-driven deep learning methods, and greater adaptability and generalization than model-based approaches. Hence, unfolding has become widely used in inverse imaging problems, such as image restoration, reconstruction, and super-resolution, but has not been sufficiently explored yet in the context of clustering. In this work, we introduce an innovative clustering architecture for hyperspectral images (HSI) by unfolding an iterative solver based on the Alternating Direction Method of Multipliers (ADMM) for sparse subspace clustering. To our knowledge, this is the first attempt to apply unfolding ADMM for computing the self-representation matrix in subspace clustering. Moreover, our approach captures well the structural characteristics of HSI data by employing the K nearest neighbors algorithm as part of a structure preservation module. Experimental evaluation of three established HSI datasets shows clearly the potential of the unfolding approach in HSI clustering and even demonstrates superior performance compared to state-of-the-art techniques.
@@ I Introduction
Subspace clustering resolves these challenges by partitioning high-dimensional data into lower-dimensional subspaces, each cluster corresponding to a distinct class. Within each subspace, data points are expressed as linear combinations of each other, forming self-representation matrices. Then a symmetric non-negative similarity matrix is obtained from the self-representation matrix. At last, spectral clustering is applied to get the cluster assignment. In subspace clustering, traditional optimizing methods such as the Alternating Direction Method of Multipliers (ADMM)[9] are applied to solve the optimization problem that yields the self-representation matrices. Traditional subspace clustering employs matrix representation to obtain shallow features for clustering, which have difficulties in handling non linearly separable data. Thus, deep clustering methods like Deep Embedded Clustering (DEC)[10] and Joint Unsupervised Learning (JULE)[11], stemming from the advancements in deep learning, have been proposed. Nevertheless, they encounter difficulties with extensive data requirements and limited interoperability.
@@ II-A Subspace clustering
The similarity matrix contains the pairwise affinity between different data points, and finally, spectral clustering is applied to get the cluster label of every data point.
@@ III Proposed method
Figure 1: The structure of the proposed method. As shown in the figure, initially, the HSI patches $x_{1},x_{2},...x_{n}$ are fed into a deep auto-encoder to extract latent representations $\mathbf{H}$, with a reconstruction loss. This latent representation is then normalized, resulting in $\tilde{\mathbf{H}}$, which is fed into an ADMM unfolding network to produce the self-representation matrix $\mathbf{C}$. To improve the quality of $\mathbf{C}$, the loss of self-representation, the loss of structure-preservation, and the sparsity loss are applied.
@@ III-A Unfolding ADMM optimization for self-representation
where $\mathbf{Z}\in\mathbb{R}^{n\times n}$ is the auxiliary variable, introduced to decompose the optimization of $\mathbf{C}$ into separate optimizations for $\mathbf{C}$ and $\mathbf{Z}$. $\mu$ and $\rho$ are used for regularizing and controlling the convergence of the optimization problem. Then we will optimize the variables one by one as follows:
@@ III-B Auto-encoder with unfolding ADMM
To leverage the spatial information in HSI data and enhance the handling of nonlinear features, we apply an auto-encoder to jointly optimize with the unfolding network. First, to ensure the latent representation retains substantial information from original data, the mean square error loss is applied:
@@
where $\tilde{\mathbf{H}}$ is the normalized latent code of original data.
@@ III-C Structure preservation module
Here, $\mathbf{A}$ represents the adjacency matrix, where $\mathbf{A}_{ij}=1$ indicates that the $i_{th}$ and $j_{th}$ data points are neighbors. $\mathbf{L}$ is the Laplacian matrix derived from $\mathbf{A}$, and $\mathbf{Tr}$ refers to the trace calculation. $\mathbf{C}_{i}$ corresponds to the $i_{th}$ row of the self-representation matrix, indicating the $i_{th}$ data point’s contribution to representing all other data points.
@@ IV-B Experiment setting and training strategy
To ensure a good initial state for the optimization process, we initialize $\mathbf{Z}$ using the 30 nearest neighbors, taking advantage of the structural prior. For $\mathbf{C}$, aiming for higher accuracy, we incorporate stricter structural information by selecting the 10 nearest neighbors. For the Salinas dataset, we use 2 ADMM iterations and a small $\rho$. For the Indian Pines and PaviaU dataset, due to more complex features, we use 3 ADMM iterations and larger $\rho$ with larger update step size. We employ three widely recognized metrics in the HSI clustering domain: Accuracy (ACC), Normalized Mutual Information (NMI), and Kappa. Higher values in these metrics correspond to better performance. We provide our code and test scripts in an online repository [^1] https://github.com/lxlscut/Unfolding-ADMM-for-Enhanced-Subspace-Clustering-of-Hyperspectral-Images. The inner structures of different datasets vary, leading to differences in the weights of various components within the loss function. The specific weights and initial $\rho$ values for each dataset are presented in Table II.

### v3

@@ Abstract
Deep subspace clustering methods are now prominent in clustering, typically using fully connected networks and a self-representation loss function. However, these methods often struggle with overfitting and lack interpretability. In this paper, we explore a clustering approach based on deep unfolding. By unfolding iterative optimization methods into neural networks, this approach offers enhanced interpretability and reliability compared to data-driven deep learning methods, and greater adaptability and generalization than model-based approaches. Unfolding techniques has become widely used in inverse imaging problems, such as image restoration, reconstruction, and super-resolution, but has not been sufficiently explored yet in the context of clustering. In this work, we introduce an innovative clustering architecture for hyperspectral images (HSI) by unfolding an iterative solver based on the Alternating Direction Method of Multipliers (ADMM) for sparse subspace clustering. To our knowledge, this is the first attempt to apply unfolding ADMM for computing the self-representation matrix in subspace clustering. Moreover, our approach captures well the structural characteristics of HSI data by employing the K nearest neighbors algorithm as part of a structure preservation module. Experimental evaluation of three established HSI datasets shows clearly the potential of the unfolding approach in HSI clustering and even demonstrates superior performance compared to state-of-the-art techniques.
@@ I Introduction
Subspace clustering resolves these challenges by partitioning high-dimensional data into lower-dimensional subspaces, each subspace corresponds to a distinct class. Within each subspace, data points are represented as linear combinations of each other, forming a self-representation matrix. Then a symmetric non-negative similarity matrix is obtained from the self-representation matrix. At last, spectral clustering is applied to get the cluster assignment. In subspace clustering, traditional optimizing methods such as the Alternating Direction Method of Multipliers (ADMM)[9] are applied to solve the optimization problem that yields the self-representation matrices. Traditional subspace clustering employs matrix representation to obtain shallow features for clustering, which have difficulties in handling non linearly separable data. Thus, deep clustering methods like Deep Embedded Clustering (DEC)[10] and Joint Unsupervised Learning (JULE)[11] have been proposed. Nevertheless, they encounter difficulties with extensive data requirements and limited interoperability.
@@ II-A Subspace clustering
The similarity matrix contains the pairwise affinity between data points, and finally, spectral clustering is applied to get the cluster label of every data point.
@@ III Proposed method
Figure 1: The structure of the proposed method. As shown in the figure, initially, the HSI patches $x_{1},x_{2},...x_{n}$ are fed into a deep auto-encoder to extract latent representations $\mathbf{H}$, with a reconstruction loss. This latent representation is then transposed and normalized, resulting in $\tilde{\mathbf{H}}$, which is fed into an ADMM unfolding network to produce the self-representation matrix $\mathbf{C}$. To improve the quality of $\mathbf{C}$, the loss of self-representation, the loss of structure-preservation, and the sparsity loss are applied.
@@ III-A Unfolding ADMM optimization for self-representation
where $\mathbf{Z}\in\mathbb{R}^{n\times n}$ is the auxiliary variable, introduced to decompose the optimization of $\mathbf{C}$ into separate optimizations for $\mathbf{C}$ and $\mathbf{Z}$. $\mu$ and $\rho$ are used for enforcing constraints and controlling the convergence of the optimization problem. Then we will optimize the variables one by one as follows:
@@ III-B Auto-encoder with unfolding ADMM
To leverage the spatial information in HSI data and enhance the handling of nonlinear features, we apply a convolutional auto-encoder to jointly optimize with the unfolding network. First, to ensure the latent representation retains substantial information from original data, the mean square error loss is applied:
@@
where $\tilde{\mathbf{H}}\in\mathbb{R}^{l\times n}$ is the normalized transpose of the latent code matrix of original data, and $l$ is the dimension of latent code, $\mathbf{C}_{i}$ is the $i_{th}$ column of $\mathbf{C}$.
@@ III-C Structure preservation module
Here, $\mathbf{A}$ represents the adjacency matrix, where $\mathbf{A}_{ij}=1$ indicates that the $i_{th}$ and $j_{th}$ data points are neighbors. $\mathbf{L}$ is the Laplacian matrix derived from $\mathbf{A}$, and $\mathbf{Tr}$ refers to the trace calculation. $\mathbf{C}_{i}$ corresponds to the $i_{th}$ column of the self-representation matrix.
@@ IV-B Experiment setting and training strategy
To ensure a good initial state for the optimization process, we initialize $\mathbf{Z}$ using the 30 nearest neighbors, taking advantage of the structural prior. For $\mathbf{C}$, aiming for higher accuracy, we incorporate stricter structural information by selecting the 10 nearest neighbors. For the Salinas dataset, we use 2 ADMM iterations and a small $\rho$. For the Indian Pines and PaviaU dataset, due to more complex features, we use 3 ADMM iterations and larger $\rho$ with larger update step size. We employ three classical metrics in the HSI clustering domain: Accuracy (ACC), Normalized Mutual Information (NMI), and Kappa. Higher values in these metrics correspond to better performance. We provide our code and test scripts in an online repository [^1] https://github.com/lxlscut/Unfolding-ADMM-for-Enhanced-Subspace-Clustering-of-Hyperspectral-Images. The inner structures of different datasets vary, leading to differences in the weights of various components within the loss function. The specific weights and initial $\rho$ values for each dataset are presented in Table II.